OpenStudy (anonymous):

IF I DONT GET ANY HELP WITH MY SEMESTER TEST I FAIL MATH AND HAVE TO REDO THE HOLE SCHOOL YEAR

OpenStudy (chosenmatt):

:(

OpenStudy (anonymous):

im planning on getting a bunch of people so we can vote the answer

OpenStudy (anonymous):

Simplify. –7 + 2(x – 3) A. x – 10 B. x – 13 C. 2x – 10 D. 2x – 13

OpenStudy (chosenmatt):

k

OpenStudy (anonymous):

theres the first one

OpenStudy (anonymous):

@moralmer000

OpenStudy (chosenmatt):

alternate form of that is 2x-13 im 100% sure

OpenStudy (anonymous):

ok, sorry i was just helping my mom by holding shamus

OpenStudy (anonymous):

Let's simplify step-by-step. −7+2(x−3) Distribute: =−7+(2)(x)+(2)(−3) =−7+2x+−6 Combine Like Terms: =−7+2x+−6 =(2x)+(−7+−6) =2x+−13 Answer: =2x−13

OpenStudy (anonymous):

Solve. m^-5=1/32 m =

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html here is the website

OpenStudy (anonymous):

how do i put that in?

OpenStudy (anonymous):

the m^-5=1/32m= becuase it wont let me

OpenStudy (anonymous):

it wouldnt le t me either

OpenStudy (anonymous):

only if you take out the ^ and m= so it would be like this m-5=1/32

OpenStudy (anonymous):

well moralmer can help right? you did yesterday and i got all 100%'s

OpenStudy (anonymous):

helloooo... @moralmer000

OpenStudy (anonymous):

earth to moralmer eath to moralmer, there seems to be no connection :( xD please i am going to fail if i dont get this right